# CVPT: Cross-Attention help Visual Prompt Tuning adapt visual task

### v1

@@ Abstract
In recent years, the rapid expansion of model sizes has led to large-scale pre-trained models demonstrating remarkable capabilities. Consequently, there has been a trend towards increasing the scale of models. However, this trend introduces significant challenges, including substantial computational costs of training and transfer to downstream tasks. To address these issues, Parameter-Efficient Fine-Tuning (PEFT) methods have been introduced. These methods optimize large-scale pre-trained models for specific tasks by fine-tuning a select group of parameters. Among these PEFT methods, adapter-based and prompt-based methods are the primary techniques. Specifically, in the field of visual fine-tuning, adapters gain prominence over prompts because of the latter’s relatively weaker performance and efficiency. Under the circumstances, we refine the widely-used Visual Prompt Tuning (VPT) method, proposing Cross Visual Prompt Tuning (CVPT). CVPT calculates cross-attention between the prompt tokens and the embedded tokens, which allows us to compute the semantic relationship between them and conduct the fine-tuning of models exactly to adapt visual tasks better. Furthermore, we introduce the weight-sharing mechanism to initialize the parameters of cross-attention, which avoids massive learnable parameters from cross-attention and enhances the representative capability of cross-attention. We conduct comprehensive testing across 25 datasets and the result indicates that CVPT significantly improves VPT’s performance and efficiency in visual tasks. For example, on the VTAB-1K benchmark, CVPT outperforms VPT over 4% in average accuracy, rivaling the advanced adapter-based methods in performance and efficiency. Our experiments confirm that prompt-based methods can achieve exceptional results in visual fine-tuning. The code is available at https://github.com/Xlgsyzp/CVPT/tree/main
@@ 1 Introduction
Increasing the scale of the models is a common method to enhance the model’s performance clip bert bart Roberta . In recent years, with the rapid development of computing devices, model sizes have significantly increased xlnet clark2020electra deberta vitg . For instance, the number of parameters in the GPT series developed by OpenAI has surged from 117 million to 1.8 trillion in just five years gpt-1 gpt-2 gpt3 . The rapidly increasing number of parameters will lead to the problem of immense computational overhead. Therefore, adapting those models to downstream tasks with the full-tuning method will incur enormous costs. To resolve this issue, the PEFT approach has been proposed lora prompt-tuning VP vl-adapter vit-adapter . PEFT adapts those large-scale pre-trained models to downstream tasks in a more efficient way by fine-tuning a subset of the models that contains much fewer parameters. Two mainstream methods within PEFT are Adapter adapter and Prompt prompt-tuning . During the training process, the Adapter inserts adapters into each transformer block and tunes those adapters, while the Prompt inserts prompt tokens into the embedded tokens to update the prompt tokens.
VPT, a prompt-based method is first introduced by Jia et al. vpt for visual fine-tuning tasks. Nevertheless, research on the adapter-based method is prominent due to its superior performance. Although some works have improved the performance of VPT dam e2vpt express , it is still challenging to match the effectiveness to that of adapter-based methods. There appears to be a consensus that prompt-based methods underperform adapter-based methods in the visual domain. But is this the case?
We conduct extensive experiments and analyses on VPT to uncover the reasons for its weaker performance compared to the Adapter. According to our experiments, we consider that the primary reason for the performance difference between VPT and adapters is that VPT’s deployment directly applies that used in NLP tasks prompt-tuning , without any adaptation to visual tasks. In NLP tasks, prompts usually contain rich semantic information that guides the fine-tuning process of the model. However, in visual tasks, prompts lack representation information. Therefore, it is necessary for VPT to use an abundant amount of prompts to fine-tune models. However, the design of VPT leads to computational inefficiency and redundancy, as well as the disruption of the self-attention between embedded tokens 3.1. As the graph follows 1, VPT shows a significant decrease in performance and an increase in costs when given a large number of prompts. Considering that, we think that VPT is unusable when given a large number of prompts.
@@
To handle the problem, we redesign VPT and introduced Cross Visual Prompt Tuning (CVPT). For the prompt tokens in CVPT, we calculate the cross-attention with the embedded tokens and add the result as residuals to the embedded tokens. This approach avoids the computational complexity of self-attention that is quadratically related to the number of prompts and allows prompts to focus on the embedded token to adapt to downstream tasks more efficiently. Additionally, by maintaining consistency in token dimensions throughout the computation process, the results of cross-attention can be directly summed with embedded tokens as residuals and do not introduce additional computational overhead for subsequent MLP. Furthermore, we share the weights of the self-attention layer with the cross-attention layer during loading checkpoints, keeping the cross-attention layer frozen alongside the self-attention layer, which eliminates the requirement for additional learned parameters for the cross-attention, and utilizes the encoded information in self-attention to help the fine-tuning of the model.
@@
We validate the effectiveness of our method on 25 datasets, the results show that the CVPT achieves a significant improvement in performance and efficiency compared to the VPT. CVPT shows an average 4% improvement in accuracy on the 19 VTAB-1K datasets, 1% on the 5 FGVC datasets, and 3% on the ADE20K dataset. Additionally, if given fewer prompt tokens, CVPT achieves a comparable performance with other advanced PEFT methods which significantly outperforms the other prompt-based methods and needs fewer learnable parameters. If a large number of prompts is allowed, our CVPT outperforms the SOTA methods on FGVC and ADE20K datasets. Besides, although a large number of prompts are inserted, it does not introduce too much extra computational overhead compared to VPT.
@@
Finally, we explore the impact of the deployment’s position and the effectiveness of the weight-sharing mechanism. The improvement on the model can be fully illustrated by the experimental results above, indicating that prompt-based methods can also rival SOTA adapter-based methods.
@@
  • We provide a detailed analysis of the application of VPT to visual tasks, and propose that its drawback can be summarised in three points which are lack of adaptation, computational inefficiency and redundancy, destruction of self-attention.
  • We propose CVPT, which introduces cross-attention and weight-sharing mechanisms, to avoid the efficiency and performance problems caused by VPT, which allows us to use more prompts to improve performance efficiently.
  • We conducted experiments on 25 datasets with different downstream tasks. The results show that our approach significantly outperforms the original VPT and other prompt-based works in terms of performance and efficiency. It is also comparable to SOTA adapter-based methods, demonstrating the usability of the prompt-based approach for visual fine-tuning.
@@ 2 Related Work
PEFT. In the era of CNN, making bigger and deeper models was an effective way to improve performance alex resnet resnext . With the rise of transformers, this trend became even more popular. The introduction of ChatGPT further cemented the goal of the community to develop larger and more powerful models. However, limited by their scale, despite their powerful performance and generality, these large models are difficult to adapt downstream tasks by using traditional paradigms (full-tuning). Consequently, NLP researchers first proposed PEFT methods. Their works demonstrate that fine-tuning just a small number of parameters in a large-scale pre-trained model can achieve nearly the same performance as full-tuning. Encouraged by the success in NLP, researchers began to apply PEFT to large-scale vision models on different visual tasks bayesian-prompt aim . After development in the past several years, the mainstream PEFT methods can be broadly categorized into adapter-based methods and Prompt-based methods.
Adapter. Jie et al. adapter proposed inserting adapters into the network to efficiently fine-tune the model. These adapters are commonly a small network that usually contains an upsampling layer and a downsampling layer. The input is multiplied with a scaling factor $\gamma$ after passing through the upsampling and downsampling layers and then the result is added as a residual to the input. The general form of adapter can be expressed as:
$X_{out}=X_{in}+\gamma(W_{up}(W_{down}(X_{in}))),$ (1)
where $X_{in}$ denotes the input of Adapter, $\gamma$ represents the scaling factor of Adapter, and $W_{up}$ and $W_{down}$ correspond to the upsampling layer and downsampling layer, respectively. Some works did some adaption to visual tasks based on Adapter, developing several variants such as AdaptFormer adaptformer , LoRA lora and RepAdapter repAdapter etc. These adapter-based methods dominate the field of visual fine-tuning.
Prompt. Prompt was originally used in the field of NLP which is added to the input text for comprehension tasks. Lester et al. prompt-tuning proposed treating the prompt as a continuous vector and fine-tuning the model by updating its gradients. Jia et al. vpt introduced this concept to visual fine-tuning for the first time, naming it VPT. As shown in Fig.3, the embedded tokens are spliced with the prompt tokens before entering each transformer block, allowing it to participate in every layer of the network within the transformer block. Before entering the next transformer block, the prompt tokens of the previous layer are discarded, and new prompt tokens are spliced with the embedded token again (VPT-Deep). This can be formulated as shown below:
@@
where the red and blue indicate learnable and frozen parameters, respectively. $P$ denotes a learnable d-dimensional vector, X is the CLS token, and E is the patched image. Although there are improved variants based on VPT, such as E2VPT e2vpt , EXPRESS express and DAM-VP dam , a performance gap remains between prompt-based and adapter-based approaches.
@@ 3 Method
3.1 Analysis of previous VPT
Firstly, we analyze VPT deeply to explore why it is not better than adapter in terms of performance and efficiency, our analysis follows three points:
Lack of adaptation to visual tasks. In NLP, each token represents an actual word with rich semantic information. Therefore, the processing of concatenating prompt tokens and embedded tokens is natural and suitable for NLP tasks. However, in visual tasks, tokens represent image patches and contain sparse semantic information compared to those in NLP. Therefore, simply splicing the prompt tokens with the embedded tokens may not provide sufficient guidance information. Additionally, visual tasks often require a deeper understanding of spatial relationships and structural features of an image, which are difficult to achieve with prompt tokens.
Computational inefficiency and redundancy. When computing self-attention, the attention between each token and all other tokens needs to be calculated. Its computational complexity is $n^{2}$, where $n$ is the number of embedded tokens. If $m$ represents the number of inserted prompt tokens, the computational complexity of self-attention in VPT can be expressed as $(n+m)^{2}$. This increases the computational overhead significantly, especially when using a larger number of prompt tokens. Additionally, we found that prompt tokens are involved in the MLP computation process, which not only adds computational overhead but also does not impact the results. Our experiments show that removing the prompt token after self-attention does not affect the results.
Destruction of self-attention between embedded tokens. After softmax, the sum of the weights of all tokens is normalized to 1. Whereas, due to the addition of the prompt tokens, the sum of the weights of the embedded tokens is reduced by the prompt tokens, which corresponds to the weakening of the representation ability of the self-attention between embedded tokens. Since the prompt token is eventually removed, this is equivalent to multiplying the self-attention result between the embedded tokens by a factor which less than one. To explore how large this effect is, we set the number of prompts to 1,5,20,50,100,150,196 respectively, and visualize the tensor after the softmax function, the results are shown in Fig.2 below.
Figure 2: Self-attention weight obtained by prompt tokens and embedded tokens. We visualize the self-attention of $cls_{token}$ and exclude itself to observe the attention of $cls_{token}$ to other tokens. And the darker the color, the larger the weight. When giving 196 prompts, the attention weight obtained by prompts is over 80%, which greatly influences the self-attention received by embedded tokens.
3.2 Cross Visual Prompt Tuning
Cross-Attention. Unlike self-attention vaswani2017attention , which computes the relationship between each element in the input sequence, cross-attention computes attention on two different sequences to process the semantic relationship between them crossvit . For example, in translation tasks, cross-attention is used to compute the attention weights between the source language sentence and the target language sentence. In our method, we introduce cross-attention to handle the semantic relationship between embedded tokens and prompt tokens, guiding the fine-tuning of the model. Specifically, the input of cross-attention consists of two parts: $X_{1}$ and $X_{2}$, in which $X_{1}\in\mathbb{R}^{n\times d_{1}}$ and $X_{2}\in\mathbb{R}^{m\times d_{2}}$. And $X_{1}$ serves as the query set and $X_{2}$ serves as the key-value set. We set $Q=X_{1}W^{Q}$ and $K=V=X_{2}W^{K}$, and then the cross-attention can be expressed as follows:

### v2

# CVPT: Cross Visual Prompt Tuning

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) has emerged to mitigate the computational demands of large-scale models. Within computer vision, adapter-based PEFT methods are often favored over prompt-based approaches like Visual Prompt Tuning (VPT) due to the latter’s performance and efficiency limitations. Our analysis reveals that VPT’s shortcomings stem from its prompt deployment strategy, which can distort the model’s inherent self-attention mechanism. To address this, we propose Cross Visual Prompt Tuning (CVPT). CVPT introduces a cross-attention module to directly model interactions between prompts and image tokens. This design decouples the prompts from the input sequence, preserving the original self-attention integrity while enabling efficient feature integration. Furthermore, we employ a weight-sharing mechanism for cross-attention initialization, which enhances representative capability without a large parameter overhead. Extensive experiments across 25 datasets show that CVPT significantly outperforms VPT. For instance, on the VTAB-1K benchmark, CVPT achieves over 4% higher average accuracy, rivaling leading adapter-based methods in both performance and efficiency. Our work confirms that prompt-based methods can achieve exceptional results in visual fine-tuning. The code is available at https://github.com/Lingyun0419/CVPT
@@ 1 Introduction
Increasing the scale of the models is a common method to enhance the model’s performance [44][11][34][35]. In recent years, with the rapid development of computing devices, model sizes have significantly increased [54][7][22][57]. For instance, the number of parameters in the GPT series developed by OpenAI has surged from 117 million to 1.8 trillion in just five years [42][43][3]. The rapidly increasing number of parameters will lead to the problem of immense computational overhead. Therefore, adapting those models to downstream tasks with the full-tuning method will incur enormous costs. To resolve this issue, the PEFT approach has been proposed [25][33][1][45][6]. PEFT adapts those large-scale pre-trained models to downstream tasks in a more efficient way by fine-tuning a subset of the models that contain much fewer parameters. Two mainstream methods within PEFT are Adapter [24] and Prompt [33]. During the training process, the Adapter inserts adapters into each transformer block and tunes those adapters, while the Prompt inserts prompt tokens into the embedded tokens to update the prompt tokens.
@@
VPT, a prompt-based method is first introduced by Jia et al. [27] for visual fine-tuning tasks. Nevertheless, research on the adapter-based method is prominent due to its superior performance. Although some works have improved the performance of VPT [26][18][9], it is still challenging to match the effectiveness to that of adapter-based methods. There appears to be a consensus that prompt-based methods underperform adapter-based methods in the visual domain. But is that really the case?
We conduct extensive experiments and analyses on VPT to uncover the reasons for its weaker performance compared to the Adapter. As a result of our investigation, we attribute this issue to the deployment of prompts in VPT. In VPT, prompts are concatenated with embedded tokens and processed together by the transformer blocks. Notably, this concatenation occurs along the token dimension, leading to the computational complexity of self-attention increasing quadratically with the number of prompts, which introduces computational redundancy. Moreover, the self-attention among embedded tokens is influenced by the prompt tokens, thereby distorting the original features 3. This implies that prompt-based methods are constrained to use only a limited number of prompts. However, due to computational inefficiencies and the limited learnable parameter count, this smaller number of prompts lacks the flexibility to adapt to various downstream tasks, resulting in a performance gap between VPT and Adapter methods.
@@
To address the issues above, it is necessary to alter the deployment of prompts to decouple prompts from the self-attention of embedded tokens. However, this also implies the need to find an alternative way to establish connections between embeddings and prompts, enabling prompts to participate in the fine-tuning process. With this in mind, we propose CVPT. We use cross-attention to capture the relationship between prompts and embedded tokens, incorporating the result as a residual term into the embedded tokens. This approach avoids the computational complexity of self-attention that is quadratically related to the number of prompts and allows prompts to focus on the embedded token to adapt to downstream tasks more efficiently. Additionally, by maintaining consistency in token dimensions throughout the computation process, the results of cross-attention can be directly summed with embedded tokens as residuals and do not introduce additional computational overhead for subsequent MLP. Furthermore, we share the weights of the self-attention layer with the cross-attention layer during loading checkpoints, keeping the cross-attention layer frozen alongside the self-attention layer, which eliminates the requirement for additional learned parameters for the cross-attention, and utilizes the encoded information in self-attention to help the fine-tuning of the model. In 1, we present the trends of accuracy and flops for CVPT and VPT with different numbers of prompts.
@@
We validate the effectiveness of our method on 25 datasets, the results show that the CVPT achieves a significant improvement in performance and efficiency compared to the VPT. CVPT shows an average 5% improvement in accuracy on the 19 VTAB-1K datasets, 1% on the 5 FGVC datasets, and 3% on the ADE20K dataset. Additionally, benefiting from the design of CVPT, it achieves superior performance compared to other prompt-based methods, even with a limited number of prompts. If a large number of prompts is allowed, CVPT outperforms the SOTA methods on out-of-distribution datasets like ADE20K. Besides, although a large number of prompts are used, it does not introduce too much extra computational overhead compared.
@@
Finally, we explore the impact of the deployment’s position and the effectiveness of the weight-sharing mechanism. The improvement on the model can be fully illustrated by the experimental results above, indicating that prompt-based methods can also rival advanced adapter-based methods.
@@
  • We provide a detailed analysis of the application of VPT to visual tasks, and propose that its drawback can be attributed to inefficiency and redundancy of computation and the destruction of self-attention, which are caused by the deployment of prompts.
  • We propose CVPT, which introduces cross-attention and weight-sharing mechanisms, to decouple prompts from the self-attention of embedded tokens, which allows prompts to integrate visual features efficiently. This makes it possible to use a large number of prompts to adapt to downstream tasks, thereby improving both performance and efficiency.
  • We conducted experiments on 25 datasets with different downstream tasks. The results show that our approach significantly outperforms the original VPT and other prompt-based works in terms of performance and efficiency. It is also comparable to advanced adapter-based methods, demonstrating the usability of the prompt-based approach for visual fine-tuning.
@@ 2 Related Work
PEFT. In the era of CNN, making bigger and deeper models was an effective way to improve performance [31][20][52]. With the rise of transformers, this trend became even more popular. ChatGPT’s introduction further cemented the community’s goal to develop larger and more powerful models. However, limited by their scale, despite their powerful performance and generality, these large models are difficult to adapt to downstream tasks by using traditional paradigms (full-tuning). Consequently, NLP researchers first proposed PEFT methods. Their works demonstrate that fine-tuning just a small number of parameters in a large-scale pre-trained model can achieve nearly the same performance as full-tuning. Encouraged by the success in NLP, researchers began to apply PEFT to large-scale vision models on different visual tasks [10][53]. After development in the past several years, the mainstream PEFT methods can be broadly categorized into adapter-based methods and Prompt-based methods. We mainly introduce prompt-based methods.
Prompt. Prompt was originally used in the field of NLP which is added to the input text for comprehension tasks. Lester et al. [33] proposed treating the prompt as a continuous vector and fine-tuning the model by updating its gradients. Jia et al. [27] introduced this concept to visual fine-tuning for the first time, naming it VPT. As shown in Fig.3, the embedded tokens are spliced with the prompt tokens before entering each transformer block, allowing it to participate in every layer of the network within the transformer block. Before entering the next transformer block, the prompt tokens of the previous layer are discarded, and new prompt tokens are spliced with the embedded token again (VPT-Deep). This can be formulated as shown below:
@@
where the red and blue indicate learnable and frozen parameters, respectively. $P$ denotes a learnable d-dimensional vector, X is the CLS token, and E is the patched image. There are improved variants based on VPT, such as E2VPT [18], EXPRESS [9] and DAM-VP [26]. In E2VPT, prompts are only combined with the key and value matrices in self-attention. However, these variants do not alter the deployment of prompts in VPT, leading to similar issues as those in VPT, and they are still unable to compete with advanced adapter-based methods.
3 Analysis of Previous VPT
The underperformance of VPT relative to adapter-based methods in the realm of PEFT has become a widely accepted notion. We aim to conduct an in-depth analysis of VPT to elucidate the root causes of its limitations. Specifically, employing a large number of prompts during fine-tuning can, in some cases, lead to significant performance improvements. In VPT-Deep, Jia et al. [27] utilized 50+ prompts on 13 out of 24 datasets to achieve optimal performance, while in VPT-Shallow, 19 datasets benefited from the use of 50+ prompts. Subsequent works [9, 26] have leveraged even larger numbers of prompts to enhance performance, with APT [2] employing over 1000 prompts. However, this increase in the number of prompts introduces the following challenges:
3.1 Computational inefficiency and redundancy
In VPT, prompts are concatenated with embedded tokens and transformer blocks process this combined input. Each token computes its attention score with every token when calculating multi-head self-attention. This results in the attention score matrix being divided into four parts: self-attention among the original tokens, self-attention among the prompts, attention from the original tokens to the prompts, and attention from the prompts to the original tokens. Since prompts are discarded at the end of each transformer block, we focus more on the embedded tokens that remain, specifically the self-attention matrix of the original tokens. We find that prompts do not directly affect the self-attention among the original tokens; instead, they can only indirectly fine-tune the embedded tokens by influencing other attention scores.
Besides, the computational complexity of self-attention is $n^{2}$, where $n$ is the number of embedded tokens. If $m$ represents the number of inserted prompt tokens, the computational complexity of self-attention in VPT can be expressed as $(n+m)^{2}$. This increases the computational overhead significantly, especially when using a larger number of prompt tokens.
Additionally, since prompt tokens are discarded at the end of each transformer block, we found that those prompt tokens not only add computational overhead but also do not impact the results. Our experiments show that removing those prompt tokens after self-attention does not affect the results. And this phenomenon has been similarly noted in previous studies [2].
3.2 Destruction of self-attention
As we analyzed above, the attention score matrix is divided into four parts. Subsequently, the attention score matrix passes through a softmax layer. Along the channel dimension, the sum of the weights for each matrix is normalized to 1. This means that as the number of prompt tokens increases, the attention weights assigned to the self-attention among the embedded tokens are progressively weakened. Therefore, the ability to represent self-attention between embedded tokens will be weakened. Since the prompt token is eventually removed, this is equivalent to multiplying the self-attention result between the embedded tokens by a factor which less than one. To explore how large this effect is, we set the number of prompts to 1,5,20,50,100,150,196 respectively, and visualize the tensor after the softmax function, the results are shown in Fig.2 below.
Figure 2: Self-attention weight obtained by prompt tokens and embedded tokens. We visualize the self-attention of $cls_{token}$ and exclude itself to observe the attention of $cls_{token}$ to other tokens. Thus, we can understand the impact of different tokens on the attention scores for the $cls_{token}$. And the darker the color, the larger the weight. When giving 196 prompts, the attention weight obtained by prompts is over 80%, which greatly influences the self-attention received by embedded tokens.
Previous solutions for excessive attention. Several prior studies have also recognized the detrimental effects caused by this phenomenon of attention convergence. Xiao et al. [51] posit that the performance collapse observed in large language models (LLMs) exceeding their maximum context length stems from attention convergence on non-informative initial tokens. This is attributed to the fact that, beyond the context length, initial tokens are excluded; however, because these tokens occupy a significant proportion of the softmax distribution, their exclusion induces a substantial shift in the distribution of attention scores, consequently leading to model failure. This observation parallels our own, as both are fundamentally rooted in the distributional issues of attention weights arising from the softmax function. They propose a mitigation strategy involving the introduction of additional learnable placeholder tokens to reduce the strong focus of other tokens on the initial tokens; this, in effect, acts as a form of smoothing for tokens that receive excessive attention.
Darcet et al. [8] note that DINOv2 exhibits suboptimal performance when utilized for feature extraction. This is attributed to the presence of a subset of tokens possessing exceptionally high norms, thereby inducing an "attention-squeezing" effect on the remaining tokens. To mitigate this, they introduce several learnable REG tokens into the sequence, serving to smooth the scores of tokens exhibiting outlier values.
Previous prompt-based method also identified the issue of prompt tokens‘ excessive attention. In VFPT [56], they demonstrate that the attention weights acquired by the original tokens in their VFPT are significantly higher compared to those in VPT. This indicates that their approach also serves to smooth the attention distribution.
3.3 Root causes of the limitations
We observe that excessive attention allocated to prompt tokens is a recurring issue in prior works. Approaches that increase the attention weights of original tokens or decrease the attention weights of prompt tokens might appear to be viable solutions. However, these methods still face the same fundamental challenges as the number of inserted prompt tokens increases. Furthermore, they do not resolve the issue of the additional computational burden, which scales quadratically with the number of prompts.
We re-examine these approaches and find that, in the scenarios described in [51, 8], the tokens exhibiting anomalous attention originate from the original sequence. As intrinsic components of the original sequence, these tokens carry substantial semantic information, playing a crucial role in contextual understanding for natural language processing tasks or feature recognition in visual tasks. However, in VPT, the anomalous token - prompts are artificially introduced. They do not inherently possess semantic meaning, serving solely as an indirect tuning factor. Their importance is thus not comparable to that of the outlier tokens observed in the aforementioned scenarios, removing them does not disrupt the network’s inherent capabilities. Based on this, we think that decoupling prompts from self-attention can fundamentally address the negative impact of prompts on self-attention.
@@ 4 Method
Having attributed the root cause of VPT’s limitations to the involvement of prompts in the self-attention process, we recognize that decoupling prompts from self-attention necessitates a new approach to establish a connection between prompt tokens and original tokens, thereby enabling their participation in fine-tuning. Consequently, the core challenge transforms into how to capture the semantic relationship between two distinct sequences.
Inspired by the works in NLP, it occurred to us to use cross-attention to establish the connection between prompt and original tokens. Therefore, we introduce CVPT (Cross Visual Prompt Tuning). Next, we review cross-attention.
Figure 3: Structure comparison of VPT and CVPT. In which blue represents frozen parameters and orange represents learnable.
4.1 Cross-Attention
Unlike self-attention [47], which computes the relationship between each element in the input sequence, cross-attention computes attention on two different sequences to process the semantic relationship between them [4]. For example, in translation tasks, cross-attention is used to compute the attention weights between the source language sentence and the target language sentence. In our method, we introduce cross-attention to handle the semantic relationship between embedded tokens and prompt tokens, guiding the fine-tuning of the model. Specifically, the input of cross-attention consists of two parts: $X_{1}$ and $X_{2}$, in which $X_{1}\in\mathbb{R}^{n\times d_{1}}$ and $X_{2}\in\mathbb{R}^{m\times d_{2}}$. And $X_{1}$ serves as the query set and $X_{2}$ serves as the key-value set. We set $Q=X_{1}W^{Q}$ and $K=V=X_{2}W^{K}$, and then the cross-attention can be expressed as follows: